# Neural-Network Holographic Model of the QCD Phase Transition under Lattice and HRG Constraints

### 26 Mar 2026 v1

@@ Abstract
Holographic QCD, HRG, Critical endpoint, Neural network
I Introduction
Ultra-relativistic heavy-ion collisions performed at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) have provided compelling experimental evidence for the creation of a new state of strongly interacting matter known as the Quark–Gluon Plasma (QGP) Adcox and others (2005); Aggarwal and others (2010); Chen and others (2024); Wang and Chen (2022, 2021); Wang et al. (2019); Xu et al. (2018); Jin et al. (2018); Liu et al. (2017); Zhu et al. (2025c); Li et al. (2024); Xie et al. (2024); He et al. (2023); Chen et al. (2025c); Xu et al. (2016). Understanding the properties of this strongly interacting QCD matter naturally leads to the investigation of the QCD phase diagram, which remains one of the central frontiers in modern nuclear physics. The transition from hadronic matter to the QGP is governed by non-perturbative dynamics that challenge traditional theoretical methods Guenther (2021); Aarts and others (2023); Bazavov and others (2014). While perturbative QCD provides accurate predictions at high energy scales, the low-energy regime—characterized by color confinement and chiral symmetry breaking—requires non-perturbative tools. Lattice QCD (LQCD) has successfully established that the transition at vanishing baryon chemical potential ($\mu_{B}=0$) is a smooth crossover Aoki et al. (2006); Borsanyi et al. (2013); Adam et al. (2025). However, extending these first-principles calculations to finite baryon density is hindered by the numerical sign problem Philipsen (2013); Aarts (2016); Nagata (2022); de Forcrand (2009), leaving the high-density regime and the potential existence of a Critical Endpoint (CEP) largely unexplored.
@@ I Introduction
Recently, machine learning (ML) techniques have been widely applied in high-energy physics Ma et al. (2023); Zhou et al. (2024); He et al. (2023); Wang and Li (2023); Pang (2024); Li et al. (2026); Shi (2026), emerging as powerful tools for tackling complex inverse problems and non-perturbative dynamics. In particular, ML provides a robust data-driven framework for extracting hidden structures from high-dimensional datasets and for constructing effective models beyond traditional analytical approaches. In the context of holographic QCD, several works have attempted to build data-driven holographic models by incorporating LQCD results, experimental measurements, or phenomenological constraints into neural-network architectures. These studies span a broad range of applications, including meson spectroscopy Hashimoto et al. (2022); Chen et al. (2025d); Zhang et al. (2026); Thomas Arun and Pal (2025), the QCD phase transition and critical phenomena Zeng et al. (2025); Cai et al. (2026); Deng et al. (2026), shear viscosity in strongly coupled plasma Yan et al. (2020), chiral condensate reconstruction Hashimoto et al. (2018a), and the heavy-quark potential Mansouri et al. (2024); Luo et al. (2025). Additional related developments can be found in Hashimoto et al. (2018b); Jeong et al. (2025); Ahn et al. (2025, 2024); Song et al. (2021); Zhu et al. (2026); Jeong et al. (2025).
@@
In our previous studies, holographic models were constructed through gradient-based parameter optimization constrained by LQCD data Chen and Huang (2024, 2025); Chen et al. (2025e). Since the Hadron Resonance Gas (HRG) model Andronic et al. (2018) provides a reliable non-perturbative description of QCD thermodynamics in the confined phase, particularly for $T\lesssim T_{c}$, it serves as a natural infrared constraint for the equation of state (EoS). Motivated by this complementarity between low- and high-temperature regimes, we develop a data-driven holographic framework that systematically incorporates high-temperature LQCD results together with low-temperature HRG constraints, allowing for a consistent reconstruction of the EoS throughout the $(T,\mu_{B})$ plane. By employing Physics-Informed Neural Network (PINN) to parameterize the bulk metric and gauge kinetic coupling, we achieve a precise description of the crossover transition without relying on rigid analytical ansätze. Our model predicts the existence of a CEP at $(T_{c},\mu_{c})=(0.089,0.922)$ GeV, connecting the smooth crossover at zero density to a first-order phase transition at high density. Additionally, we utilize symbolic regression to derive analytic expressions for the reconstructed background functions, offering a transparent holographic dual for QCD matter. The paper is organized as follows: Section II outlines the EMD formalism and the neural network training methodology. Section III presents the numerical results, including the thermodynamic variables, the phase diagram, and the symbolic regression analysis. Finally, a summary is provided in Section IV.
II Theoretical Descriptions
II.1 Einstein-Maxwell-Dilaton Framework
@@
$S=\frac{1}{16\pi G_{5}}\int d^{5}x\,\sqrt{-g}\bigg[$ $R-\frac{f(\phi)}{4}F^{2}$
$-\frac{1}{2}\,\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)\bigg],$ (1)
@@
II.2 Neural Network Reconstruction
@@
This allows the model to learn the non-trivial density dependence implicitly contained in the lattice baryon susceptibility $\chi_{2}^{B}$, without restricting the solution space to a specific analytical class. Once $A(z)$ and $f(z)$ are fixed by the networks, the remaining background functions $g(z)$ and $A_{t}(z)$ are derived by solving the Einstein-Maxwell equations.
@@
$g(z)$ $=1+\frac{1}{\int_{0}^{z_{h}}y^{3}e^{-3A}dy}\left[-\int_{0}^{z}y^{3}e^{-3A}dy\right$.
$\left.+\left(\frac{\mu}{\int_{0}^{z_{h}}\frac{y}{e^{A}f}dy}\right)^{2}\hat{G}\right$],
@@
$V(z)$ $=-3z^{2}ge^{-2A}\bigg[A^{\prime\prime}+A^{\prime}\left(3A^{\prime}-\frac{6}{z}+\frac{3g^{\prime}}{2g}\right)$
$\quad-\frac{1}{z}\left(-\frac{4}{z}+\frac{3g^{\prime}}{2g}\right)+\frac{g^{\prime\prime}}{6g}\bigg],$
@@
II.3 Thermodynamics and Conserved Quantities
@@
$T=$ $\frac{1}{4\pi}\frac{z_{h}^{3}e^{-3A\left(z_{h}\right)}}{\int_{0}^{z_{h}}y^{3}e^{-3A(z)}dy}\left[1-\left(\frac{\mu}{\int_{0}^{z_{h}}\frac{y}{e^{A}f}dy}\right)^{2}\right$. (7)
$\quad\left(\int_{0}^{z_{h}}y^{3}e^{-3A}dy\cdot\int_{0}^{z_{h}}\frac{x}{e^{A}f}dx\right$.
$\quad\left.\left.-\int_{0}^{z_{h}}y^{3}e^{-3A}dy\int_{0}^{y}\frac{x}{e^{A}f}dx\right)\right],$
@@
Given the entropy and density $\rho$, the free energy and pressure follow from integration of the appropriate thermodynamic identity Critelli et al. (2017); Chen and Huang (2025).
@@
The free energy is normalized to zero at $(T,\mu)=(0,0)$. With this convention, the energy density can be expressed as
@@ II.4 Training Strategy
To address the scarcity of lattice data in the low-temperature regime, we augment the training dataset with HRG model predictions for $T<0.13$ GeV. This hybrid dataset ensures that the neural network learns the correct hadronic degrees of freedom, thereby determining more precise thermodynamic properties.
III Results and discussion
A fundamental challenge in bottom-up holographic QCD is the construction of a bulk geometry that simultaneously captures the physics of the deconfined QGP and the confined hadronic phase within a unified framework. Traditional models often rely on hand-crafted analytical ansätze for the warp factor $A(z)$ or dilaton potential $V(\phi)$, which inherently introduces human-imposed bias and often fails to reproduce the hadronic limit at low temperatures ($T\leq 0.13$ GeV).
In this work, we implement a data-to-holography pipeline that solves the inverse problem by treating $A(z)$ and the gauge-dilaton coupling $f(z)$ as deep-learning-optimized functional profiles. To anchor the model across the entire phase space, we adopt a hybrid training strategy: (2+1)-flavor LQCD data Bazavov and others (2014); Bellwied et al. (2015) is used for the high-temperature regime. For the low-temperature region, where lattice calculations encounter computational bottlenecks, the model is constrained by predictions from the HRG model, these predictions are generated using the Thermal-FIST package Vovchenko and Stoecker (2019), a C++-based software tool specifically designed for studying heavy-ion collisions and hadronic equations of state.
III.1 Thermodynamic Validation at $\mu_{B}=0$
At vanishing chemical potential, our machine-learning augmented EMD model demonstrates quantitative agreement with first-principles thermodynamics. As shown in Fig.1, where we present the temperature dependence of the thermodynamic EoS at vanishing baryon chemical potential, the normalized pressure $p$, energy density $\epsilon$, and entropy density $s$ align closely with lattice benchmarks across the range $0.13\text{ GeV}\leq T\leq 0.60\text{ GeV}$.
Figure 1: Comparison of NNHM predictions with LQCD results for thermodynamic quantities at vanishing chemical potential. The blue solid lines represent the NNHM results, while the red points with error bars denote the (2+1)-flavor LQCD data from Ref. Bazavov and others (2014); Bellwied et al. (2015). For reference, the green solid lines show the HRG model calculations, and the purple dashed lines indicate the SB limits. The plotted quantities are: entropy density $s$, energy density $\epsilon$, pressure $P$, trace anomaly $\epsilon-3P$, second-order baryon susceptibility $\chi_{2}^{B}$, squared speed of sound $C_{s}^{2}$, specific heat $C_{V}$, and free energy $F$, all appropriately scaled by powers of temperature $T$.
@@ III.1 Thermodynamic Validation at $\mu_{B}=0$
III.2 Finite Density Extrapolation and Susceptibility Response
@@
III.3 Localization of the CEP and Phase Structure
@@
This topological structure is qualitatively consistent with predictions from functional approaches, including the Functional Renormalization Group (FRG) Fu et al. (2020); Zhang et al. (2017); Osman et al. (2025); Roth et al. ; Murgana et al. (2024); Isserstedt et al. (2019) and Dyson–Schwinger Equations (DSE) Gao and Liu (2016); Qin et al. (2011); Shi et al. (2014); Fischer et al. (2014). Compared with previous holographic studies, the CEP obtained in our model appears at a larger chemical potential due to the inclusion of HRG constraints. To illustrate this effect, several representative results are shown in Fig. 3. The brown hexagon denotes a potential-reconstruction EMD model, which lies close to the $V(\phi)$ EMD model represented by the orange triangle Yang et al. (2026). The purple square corresponds to the merged EoS model incorporating both LQCD and HRG data Yang et al. (2026). These comparisons demonstrate that the inclusion of HRG data significantly shifts the CEP location toward larger baryon chemical potential within the holographic framework.
III.4 Analytic Reconstruction via Symbolic Regression
@@
$A(z)=z\left[\left(a+\frac{b}{c-\left(z+\frac{d}{z}\right)}\right)(z+e)\right]$ (15)
l
$f(z)=f_{0}\exp\left[f+\frac{g}{h\left(i+\frac{1}{j-z^{2}}\right)^{2}+k}\right]$ (16)
@@
III.5 Three-Dimensional Structure of the QCD Equation of State
@@
III.6 Data-driven verification
@@
Quantitatively, the CEP (yellow diamond) extracted from data generated by the potential-reconstruction EMD model is located at $(T_{c},\mu_{c})=(0.116,0.656)\mathrm{GeV}$, which is close to the value obtained in the original model, $(T_{c},\mu_{c})=(0.110,0.678)\mathrm{GeV}$, reported in Ref. Chen et al. (2025e), within the expected numerical uncertainties. This agreement supports the reliability of our reconstruction pipeline and indicates that the CEP prediction of the NNHM is genuinely data-driven.  [^1] The code can be found at: https://github.com/De-Xing-Zhu/
IV Summary

### 30 Mar 2026 v2

# Wuhan 430079, China

@@ Abstract
arxiv: xxxx
1 Introduction
Ultra-relativistic heavy-ion collisions performed at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) have provided compelling experimental evidence for the creation of a new state of strongly interacting matter known as the Quark–Gluon Plasma (QGP) Abelev and others (2012); Chatrchyan and others (2014); Adam and others (2015); Khachatryan and others (2017); Acharya and others (2017); Sirunyan and others (2017, 2020); Acharya and others (2021); Muller et al. (2012); Busza et al. (2018); Adcox and others (2005); Aggarwal and others (2010); Chen and others (2024); Wang and Chen (2022, 2021); Wang et al. (2019); Xu et al. (2018); Jin et al. (2018); Liu et al. (2017); Zhu et al. (2025c); Li et al. (2024); Xie et al. (2024); He et al. (2023); Chen et al. (2025c); Xu et al. (2016). Understanding the properties of this strongly interacting QCD matter naturally leads to the investigation of the QCD phase diagram, which remains one of the central frontiers in modern nuclear physics. The transition from hadronic matter to the QGP is governed by non-perturbative dynamics that challenge traditional theoretical methods Guenther (2021); Aarts and others (2023); Bazavov and others (2014). While perturbative QCD provides accurate predictions at high energy scales, the low-energy regime—characterized by color confinement and chiral symmetry breaking—requires non-perturbative tools. Lattice QCD (LQCD) has successfully established that the transition at vanishing baryon chemical potential ($\mu_{B}=0$) is a smooth crossover Aoki et al. (2006); Borsanyi et al. (2013); Adam et al. (2025). However, extending these first-principles calculations to finite baryon density is hindered by the numerical sign problem Philipsen (2013); Aarts (2016); Nagata (2022); de Forcrand (2009), leaving the high-density regime and the potential existence of a Critical Endpoint (CEP) largely unexplored.
@@ 1 Introduction
Recently, machine learning (ML) techniques have been widely applied in high-energy physics Ma et al. (2023); Zhou et al. (2024); He et al. (2023); Wang and Li (2023); Pang (2024); Li et al. (2026); Shi (2026), emerging as powerful tools for tackling complex inverse problems and non-perturbative dynamics. In particular, ML provides a robust data-driven framework for extracting hidden structures from high-dimensional datasets and for constructing effective models beyond traditional analytical approaches. In the context of holographic QCD, several works have attempted to build data-driven holographic models by incorporating LQCD results, experimental measurements, or phenomenological constraints into neural-network architectures. These studies span a broad range of applications, including meson spectroscopy Hashimoto et al. (2022); Chen et al. (2025d); Zhang et al. (2026); Thomas Arun and Pal (2025), the QCD phase transition and critical phenomena Zeng et al. (2025); Cai et al. (2026); Deng et al. (2026), shear viscosity in strongly coupled plasma Yan et al. (2020), chiral condensate reconstruction Hashimoto et al. (2018a), and the heavy-quark potential Mansouri et al. (2024); Luo et al. (2025). Additional related developments can be found in Hashimoto et al. (2018b); Ahn et al. (2025, 2024); Song et al. (2021); Zhu et al. (2026); Jeong et al. (2025).
@@
In our previous studies, holographic models were constructed through gradient-based parameter optimization constrained by LQCD data Chen and Huang (2024, 2025); Chen et al. (2025e). Since the Hadron Resonance Gas (HRG) model Andronic et al. (2018) provides a reliable non-perturbative description of QCD thermodynamics in the confined phase, particularly for $T\lesssim T_{c}$, it serves as a natural infrared constraint for the equation of state (EoS). Motivated by this complementarity between low- and high-temperature regimes, we develop a data-driven holographic framework that systematically incorporates high-temperature LQCD results together with low-temperature HRG constraints, allowing for a consistent reconstruction of the EoS throughout the $(T,\mu_{B})$ plane. By employing Physics-Informed Neural Network (PINN) to parameterize the bulk metric and gauge kinetic coupling, we achieve a precise description of the crossover transition without relying on rigid analytical ansätze. Our model predicts the existence of a CEP at $(T_{c},\mu_{c})=(0.089,0.922)$ GeV, connecting the smooth crossover at zero density to a first-order phase transition at high density. Additionally, we utilize symbolic regression to derive analytic expressions for the reconstructed background functions, offering a transparent holographic dual for QCD matter. The paper is organized as follows: Section 2 outlines the EMD formalism and the neural network training methodology. Section 3 presents the numerical results, including the thermodynamic variables, the phase diagram, and the symbolic regression analysis. Finally, a summary is provided in Section 4.
2 Theoretical Descriptions
2.1 Einstein-Maxwell-Dilaton Framework
@@
$S=\frac{1}{16\pi G_{5}}\int d^{5}x\,\sqrt{-g}\bigg[R-\frac{f(\phi)}{4}F^{2}-\frac{1}{2}\,\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)\bigg],$ (1)
@@
2.2 Neural Network Reconstruction
@@
This allows the model to learn the non-trivial density dependence implicitly contained in the lattice baryon susceptibility $\chi_{2}^{B}$, without restricting the solution space to a specific analytical class. Once $A(z)$ and $f(z)$ are fixed by the networks, the remaining background functions $g(z)$ and $A_{t}(z)$ are derived by solving the Einstein-Maxwell equations:
@@
$g(z)=1+\frac{1}{\int_{0}^{z_{h}}y^{3}e^{-3A}dy}\bigg[-\int_{0}^{z}y^{3}e^{-3A}dy+\left(\frac{\mu}{\int_{0}^{z_{h}}\frac{y}{e^{A}f}dy}\right)^{2}\hat{G}\bigg$],
@@
$V(z)=-3z^{2}ge^{-2A}\bigg[A^{\prime\prime}+A^{\prime}\left(3A^{\prime}-\frac{6}{z}+\frac{3g^{\prime}}{2g}\right)-\frac{1}{z}\left(-\frac{4}{z}+\frac{3g^{\prime}}{2g}\right)+\frac{g^{\prime\prime}}{6g}\bigg],$
@@
2.3 Thermodynamics and Conserved Quantities
@@
$T=$ $\frac{1}{4\pi}\frac{z_{h}^{3}e^{-3A\left(z_{h}\right)}}{\int_{0}^{z_{h}}y^{3}e^{-3A(z)}dy}\times\bigg[1-\left(\frac{\mu}{\int_{0}^{z_{h}}\frac{y}{e^{A}f}dy}\right)^{2}$ (7)
$\qquad\times\left(\int_{0}^{z_{h}}y^{3}e^{-3A}dy\cdot\int_{0}^{z_{h}}\frac{x}{e^{A}f}dx-\int_{0}^{z_{h}}y^{3}e^{-3A}dy\int_{0}^{y}\frac{x}{e^{A}f}dx\right)\bigg],$
@@
Given the $s$ and $\rho$, the free energy $F$ and pressure $P$ follow from integration of the appropriate thermodynamic identity Critelli et al. (2017); Chen and Huang (2025).
@@
The $F$ is normalized to zero at $(T,\mu)=(0,0)$. With this convention, the energy density can be expressed as
@@ 2.4 Training Strategy
To address the scarcity of lattice data in the low-temperature regime, we augment the training dataset with HRG model predictions for $T<0.13$ GeV. This hybrid dataset ensures that the neural network learns the correct hadronic degrees of freedom, thereby determining more precise thermodynamic properties. For this study, $G_{5}$ = 4.17069 was obtained through neural network training.  [^1] The code can be found at: https://github.com/De-Xing-Zhu/
3 Results and discussion
A fundamental challenge in bottom-up holographic QCD is the construction of a bulk geometry that simultaneously captures the physics of the deconfined QGP and the confined hadronic phase within a unified framework. Traditional models often rely on hand-crafted analytical ansätze for the warp factor $A(z)$ or dilaton potential $V(\phi)$, which inherently introduces human-imposed bias and often fails to reproduce the hadronic limit at low temperatures ($T<0.13$ GeV).
In this work, we implement a data-to-holography pipeline that solves the inverse problem by treating $A(z)$ and the gauge-dilaton coupling $f(z)$ as deep-learning-optimized functional profiles. To anchor the model across the entire phase space, we adopt a hybrid training strategy: (2+1)–flavor LQCD data Bazavov and others (2014); Bellwied et al. (2015) is used for the high-temperature regime. For the low-temperature region, where lattice calculations encounter computational bottlenecks, the model is constrained by predictions from the HRG model, these predictions are generated using the Thermal-FIST package Vovchenko and Stoecker (2019), a C++ based software tool specifically designed for studying heavy-ion collisions and hadronic equations of state.
3.1 Thermodynamic Validation at $\mu_{B}=0$
At vanishing chemical potential, our machine-learning augmented EMD model demonstrates quantitative agreement with first-principles thermodynamics. As shown in Fig.1, where we present the temperature dependence of the thermodynamic EoS at vanishing baryon chemical potential, the normalized pressure $P$, energy density $\epsilon$, and entropy density $s$ align closely with lattice benchmarks across the range $0.13\text{ GeV}\leqslant T\leqslant 0.60\text{ GeV}$.
Figure 1: Comparison of NNHM predictions with LQCD results for thermodynamic quantities at vanishing chemical potential. The blue solid lines represent the NNHM results, while the red points with error bars denote the (2+1)–flavor LQCD data from Ref. Bazavov and others (2014); Bellwied et al. (2015). For reference, the green solid lines show the HRG model calculations, and the purple dashed lines indicate the SB limits. The plotted quantities are: entropy density $s$, energy density $\epsilon$, pressure $P$, trace anomaly $\epsilon-3P$, second-order baryon susceptibility $\chi_{2}^{B}$, squared speed of sound $C_{s}^{2}$, specific heat $C_{V}$, and free energy $F$, all appropriately scaled by powers of temperature $T$.
@@ 3.1 Thermodynamic Validation at $\mu_{B}=0$
3.2 Finite Density Extrapolation and Susceptibility Response
@@
3.3 Localization of the CEP and Phase Structure
@@
This topological structure is qualitatively consistent with predictions from functional approaches, including the Functional Renormalization Group (FRG) Fu et al. (2020); Zhang et al. (2017); Osman et al. (2025); Roth et al. ; Murgana et al. (2024); Isserstedt et al. (2019) and the Dyson-Schwinger Equations (DSE) Gao and Liu (2016); Qin et al. (2011); Shi et al. (2014); Fischer et al. (2014). Compared with previous holographic studies, the CEP obtained in our model appears at a larger chemical potential due to the inclusion of HRG constraints. To illustrate this effect, several representative results are shown in Fig. 3. The brown hexagon denotes a potential-reconstruction EMD model, which lies close to the $V(\phi)$ EMD model represented by the orange triangle Yang et al. (2026). The purple square corresponds to the merged EoS model incorporating both LQCD and HRG data Yang et al. (2026). These comparisons demonstrate that the inclusion of HRG data significantly shifts the CEP location toward larger baryon chemical potential within the holographic framework.
3.4 Analytic Reconstruction via Symbolic Regression
@@
$A(z)=z\left[\left(a+\frac{b}{c-\left(z+\frac{d}{z}\right)}\right)(z+e)\right],$ (15)
$f(z)=f_{0}\exp\left[f+\frac{g}{h\left(i+\frac{1}{j-z^{2}}\right)^{2}+k}\right].$ (16)
@@
3.5 Three-Dimensional Structure of the QCD Equation of State
@@
3.6 Data-driven verification
@@
Quantitatively, the CEP (yellow diamond) extracted from data generated by the potential-reconstruction EMD model is located at $(T_{c},\mu_{c})=(0.116,0.656)\mathrm{GeV}$, which is close to the value obtained in the original model, $(T_{c},\mu_{c})=(0.110,0.678)\mathrm{GeV}$, reported in Ref. Chen et al. (2025e), within the expected numerical uncertainties. This agreement supports the reliability of our reconstruction pipeline and indicates that the CEP prediction of the NNHM is genuinely data-driven.
4 Summary
@@
Acknowledgements.
We would like to thank Yan-Qing Zhao and Jia-Hao Wang for helpful discussions. This work is supported in part National Key Research and Development Program of China under Contract No. 2022YFA1604900 and by the National Natural Science Foundation of China (NSFC) Grant No. 12405154, No.92570117, No.12435009 and No. 12275104, Shenzhen Peacock fund under No. 2023TC0179, Ministry of Science and Technology of China under Grant No. 2024YFA1611004, the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR).